OpenStudy (anonymous):

how do i write an equation in standard form for a line passing through A(-2,3) and b(3,4)?

OpenStudy (aravindg):

use two point form

OpenStudy (anonymous):

which is what?

OpenStudy (aravindg):

\[y-y_1=(x-x_1)\dfrac{(y_2-y_1)}{(x_2-x_1)}\]

OpenStudy (aravindg):

got it ? :)

OpenStudy (anonymous):

i think maybe. one sec ima try it

OpenStudy (anonymous):

i dont get how to use it all i at y-3=(x+2)1/5

OpenStudy (anonymous):

im* at

OpenStudy (anonymous):

you are good at that step

OpenStudy (anonymous):

im not sure

OpenStudy (anonymous):

you have \[y-3=\frac{1}{5}(x+2)\] now you can multiply both sided by 5 and get \[5y-15=x+2\]

OpenStudy (anonymous):

then maybe \[x-5y=-17\]

OpenStudy (anonymous):

i think that is 'standard form"

OpenStudy (anonymous):

i know stanard form Ax+By=C

OpenStudy (anonymous):

ok then we have it

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw